Firm l produces x and Firm 2 produces y. Each firm is a price taker. The price of x is $2, and the price of y is $3.
The cost function of x is given by
x^÷100 and the cost function of y is given by y^÷100_x .How much subsidy per unit of production in x will induce a socially efficient supply?
A $0.75
B. $1.00
C. $1.22
D. $1.50​
1
Expert's answer
2021-09-27T10:49:22-0400

Solution:

The correct answer is D. $1.50


Socially efficient supply is where MSB = MSC

Firm 1 MSB (X) = 2

Firm 2 MSB (Y) = 3

Firm 1 MSC ="\frac{\partial TC} {\partial X} = \frac{X} {50}"


Firm 2 MSC = "\frac{\partial TC} {\partial Y} = \frac{Y} {50}"


Firm 1 socially efficient supply: MSB = MSC

= 2 = "\frac{X} {50}"

X = 100

Firm 2 socially efficient supply: MSB = MSC

= 3 = "\frac{Y} {50}"

Y = 150

Subsidy per unit of production in x that will induce a socially efficient supply ="\frac{Y} {X} = \frac{150} {100} = 1.50"